Can you pls solve this problem
Write an equation for the line described
a: y-intercept -3 and the same slope as the graph 2x+4y=5
b: Same slope as the graph of y-2x=3 and the same y-intercept as the graph of 3y+2x=6

--> I'm expecting for you're imediate reply for this assignment...
Thank you!

Answer by Earlsdon(6294)

Well...I don't know whether or not the timing of this reply meets your stringent expectations, but here it is anyway.
a) The equation of the line whose slope is the same as that of the line 2x + 4y = 5.
To find the slope, put the equation into the slope-intercept form: y = mx + b and then you can simply read the slope directly.
2x + 4y = 5 Subtract 2x from both sides.
4y = -2x + 5 Divide both sides by 4
y = (-2/4)x + 5/4
So, the slope, m, is -2/4 or -1/2
The y-intercept is given as -3
The required equation is: y = (-1/2)x - 3
b) For this problem, you need to find the slope from the first given equation (y-2x=3) and the y-intercept from the second given equation (3y+2=6).
Again, using the form: y = mx + b
The first equation becomes: y = 2x + 3 You get this by adding 2x to both sides.
The second equation becomes: y = (-2/3)x + 2 You get this by subtracting 2x from both sides and dividing both sides by 3.
Now you have the slope from the first equation: m = 2
and the y-intercept from the second equation: b = 2
The required equation is: y = 2x + 2